OpenStudy (anonymous):

Solve the absolute value equation |2n + 4| = 8. One solution is n = 2, what is the other solution?

OpenStudy (anonymous):

I got -6

OpenStudy (anonymous):

\[| x|= n\] can be rewritten into 2 equations x=n & x=-n

Directrix (directrix):

@trini_00x Please review the OpenStudy Code of Conduct to avoid violating it. Thanks. OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

|2 x -6 + 4| = 8 2 x -6 = -12 -12 + 4 = 8 if you get the absolute value

OpenStudy (anonymous):

HAPPY NOW @Directrix ??

OpenStudy (anonymous):

Thank you @trini_00x

OpenStudy (anonymous):

Your welcome :D